OpenStudy (anonymous):

how would The absolute graph of l7x-2l=2 look like?

OpenStudy (anonymous):

\[ 7x -2 =2\\ 7x =4\\ x= \frac 4 7\\ also\\ 7x-2 = -2\\ 7x =0\\ x=0 \]